OpenStudy (anonymous):

prove d/dx (square root of u) = 1/2*square root of u * du/dx

Parth (parthkohli):

Chain rule.

OpenStudy (anonymous):

can u help pls..

Parth (parthkohli):

\[{d \over dx} \sqrt{u} \implies {d \over dx} u^{1 \over 2} \implies {1 \over 2} \times \sqrt{u} \times {du \over dx}\]

Parth (parthkohli):

That is chain rule!

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

help me pls..

OpenStudy (zzr0ck3r):

hmm not sure this is a proof.

OpenStudy (zzr0ck3r):

what class is this for?

OpenStudy (anonymous):

differential calculus.. it is a chain rule